You would like to save up $1,000,000 in 45 years by making equal end of the year deposits into a tax-deferred account paying 6% annually. What must your annual deposit equal to achieve this goal?
Answer #1

Desired Sum = $1,000,000
Interest Rate = 6%
Time Period = 45 years

Let annual deposit be $x

$x*1.06^44 + $x*1.06^43 + … + $x*1.06 + $x = $1,000,000
$x * (1.06^45 - 1) / 0.06 = $1,000,000
$x * 212.743514 = $1,000,000
$x = $4,700.50

Annual Deposit = $4,700.50